OpenStudy (anonymous):

Which pair of ratios is proportional? A. 6/21 and 2/42 B.25/100 and 50/150 C.18/24 and 6/12 D.7/14 and 20/40

OpenStudy (texaschic101):

7/14 = 1/2 20/40 = 1/2